# Topological Kleene Field Theories: A new model of computation

### v1

@@ 2.2. Partial Recursive Functions
$f_{M}(x)=\left\{\begin{array}[]{ll}y&\textrm{if $M$ halts when $x$ is used as initial value and returns $y$},\\ \bullet&\textrm{if $M$ does not halt when $x$ is used as initial value}.\end{array}\right.$
@@
Remark 2.2.
In the previous description, the contents of the tape to the left of the head in the initial and final steps of the computation are disregarded, and this space is used as an “auxiliary” memory for the machine. This choice allows us, in particular, to implement “injective” Turing machines, in the sense that the content of the whole tape is different at the end of the computation for different inputs.
Theorem 2.3 ([22]).
The function $f_{M}:\mathbb{N}\dashrightarrow\mathbb{N}$ implemented by a binary Turing machine $M$ is a partial recursive function. Reciprocally, every partial recursive function can be implemented via a Turing machine.
Notice that the correspondence of the previous theorem is not bijective: Different Turing machines can implement the same partial recursive function.
@@ 2.3. Turing Machines as finite state machines
$\begin{array}[]{lclclclclcl}\delta(q_{0},0)&=&(q_{0},0,+1),&&\delta(q_{0},1)&=&(q_{1},0,+1),&&\delta(q_{0},\square)&=&(q_{2},1,-1),\\ \delta(q_{1},0)&=&(q_{1},0,+1),&&\delta(q_{1},1)&=&(q_{0},0,+1),&&\delta(q_{1},\square)&=&(q_{2},0,-1).\\ \end{array}$
This Turing machine is designed to read a string of 0’s and 1’s and to return $1$ if it finds an even number of 1’s and 00 otherwise. According to the previous description, the associated finite state machine is the one depicted in fig. 3. The halting states are highlighted with a double border, and the initial state is marked with an incoming edge from no node.
$q_{2}$$q_{0}$$q_{1}$$(\square,1,-1)\;$$(1,0,+1)$$(1,0,+1)$$\;(\square,0,-1)$$(0,0,+1)$$(0,0,-1)$
@@ 3.2. Cantor-based encodings
The standard ternary Cantor set is the set of real numbers $x\in\mathbb{R}$ of the form $x=\sum_{i=1}^{\infty}\epsilon_{i}3^{-i}$, with $\epsilon_{i}=0$ or $2$. This Cantor set can be used to encode natural numbers (through their binary expansion) in the real line or, even more interestingly, paths in a binary tree ($\epsilon_{i}=0$ means that we turn left in the $i$-th branch, and $\epsilon_{i}=2$ that we turn right). Due to these interesting properties, the study of its group of automorphisms is a very active area, see for instance [18, 25, 33].
Using these ideas, we can define a useful encoding of finite binary sequences with symbols from the alphabet $\mathcal{A}=\{0,1,\square\}$ into the closed interval $I=[-1,1]\subseteq\mathbb{R}$. For this purpose, let $t=\{t_{n}\}$ be finite sequence with $t_{n}\in\mathcal{A}$ for $n\geq 0$ and $t_{m}=\square$ for $m$ large enough. We encode $t$ in $I$ as the point
@@
For any $n\geq 2$, the tuple $(\mathbb{D}^{n},X,\iota)$ is a marked dynamical germ, where $\mathbb{D}^{n}$ is the standard disc of radius $2$, $\iota:I^{n}\hookrightarrow\mathbb{D}^{n}$ the usual inclusion and $X_{p}=(0,-1)$ for all $p\in\mathbb{D}^{n}$ is the normal vector field. We shall denote this marked dynamical germ just by $\mathbb{D}^{n}$, or just $\mathbb{D}$ when the dimension is clear from the context, and call it the standard disc.
@@ 3.3. Partial functions from dynamical bordisms
We may have $\psi_{(W,X)}(c)=\bullet$ for two different reasons: Either the flow starting at $x$ does not hit the outgoing boundary $M^{\prime}$ or it hits $M^{\prime}$ but not in the image of the Cantor set, as represented in Figure 4.
Figure 4. Different possibilities ocurring in the definition of $\psi_{(W,X)}:C^{n}\dashrightarrow C^{n}$.
@@
The clumsy condition that the line $\{n\}\times\kappa(\mathbb{N}^{n-1})$ must be mapped to a line $\{n^{\prime}\}\times\kappa(\mathbb{N}^{n-1})$ should be understood as the fact that the last $n-1$ components of the initial point are just “auxiliary data” that do not affect the final result of the calculation. The need for these auxiliary data is very clear geometrically. Indeed, since $\psi_{(W,X)}$ is given as the hitting map of a smooth flow, it must be injective. This would imply that the partial functions captured by $Z_{0}$ must be injective. In order to avoid this problem, we use the vertical direction to accumulate auxiliary data that allows separating points with the same images.
In the general case of a dynamical bordism $(W,X):(M,X,\bm{\iota})\to(M^{\prime},X^{\prime},\bm{\iota}^{\prime})$ between non-connected objects, decompose $M=M_{1}\sqcup\ldots\sqcup M_{m}$ and $M^{\prime}=M_{1}^{\prime}\sqcup\ldots\sqcup M^{\prime}_{m^{\prime}}$ into their connected components so that we have restrictions $\iota_{j}:I^{n}\hookrightarrow M_{j}$ and $\iota_{j}^{\prime}:I^{n}\hookrightarrow M_{j}^{\prime}$. Working as above, for any $1\leq j\leq n$, we define the auxiliary partial maps
$\psi_{(W,X),j}:\mathbb{N}^{n}\dashrightarrow(\mathbb{N}^{n})^{\sqcup m^{\prime}},\qquad\psi_{(W,X),j}(c)=\left\{\begin{array}[]{ll}(\iota_{k}^{\prime}\circ\kappa)^{-1}(x^{\prime})&\textrm{if $x^{\prime}\in\textrm{Img}(\iota_{k}^{\prime}\circ\kappa)$},\\ \bullet&\textrm{otherwise}.\end{array}\right..$
Here $x^{\prime}$ is the intersection point of $\iota_{j}(\kappa(c))\in M_{j}\subseteq W$ with $M^{\prime}$. Putting these maps together, we get a partial map
$\psi_{(W,X)}=\sqcup_{j}\psi_{(W,X),j}:(\mathbb{N}^{n})^{\sqcup m}\dashrightarrow(\mathbb{N}^{n})^{\sqcup m^{\prime}}$
that induces as in the previous case a partial function $Z_{0}(W,X):\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$. We have thus proven the following result.
@@
At this point, an important question arises: What partial functions are representable by a clean dynamical bordism? In other words, given a partial function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, does there exist a clean dynamical bordism $W_{f}$ such that $Z_{0}(W_{f})=f$? In this direction, the following theorem establishes the first part of Theorem 1.1. Its proof can be checked in Appendix A.
@@
In general, the topology of the bordism representing a partial recursive function $f:\mathbb{N}\to\mathbb{N}$ will not be trivial, capturing the complexity of the function. In particular, the function $f$ can be represented by a basic morphism if and only if $f$ extends to a diffeomorphism of the disc, when seen $\mathbb{N}\hookrightarrow\mathbb{D}$ via $\kappa$ (recall that $\mathbb{D}$ is the standard disc of radius $2$). However, not all the partial recursive functions can be extended, such as the function that reverses the binary expansion of a natural number, i.e., $f\left(\sum_{i=0}^{n}a_{i}2^{i}\right)=\sum_{i=0}^{n}a_{n-i}2^{i}$. Indeed, the points in $\mathbb{D}$ associated to the sequences $x_{n}=2^{n}$ and $x^{\prime}_{n}=2^{n}+2^{n-1}$ converge to the same limit $(-1,0)\in\mathbb{D}$, but $f(x_{n})=1$ and $f(x_{n}^{\prime})=3$ for all $n$.
@@
Recall that, by definition, the vector field of a dynamical bordism is transverse to the incoming and outgoing boundaries of the bordism. This condition is crucial in Theorem 3.9. If we drop it, it is very simple to construct a vector field on a topologically trivial bordism representing any function. Indeed, if $A\subseteq M$ is any discrete subset of a compact manifold $M$ (possibly with boundary), given any function $f:A\to A$, we can easily construct a degenerate non-transverse dynamical bordism representing it. To do so, in $M\times[0,1]$, choose pairwise disjoint paths $\gamma_{a}$ connecting $(a,0)$ with $(f(a),1)$ for all $a\in A$. Since $A$ is discrete, these paths $\gamma_{a}$ can be thickened into pairwise disjoint closed tubes $T_{a}\subseteq M\times[0,1]$. We endow each of these tubes $T_{a}$ with a flow-box vector field $X_{a}$ in such a way that the flow starting at $(a,0)$ hits $(f(a),1)$ and such that $X_{a}$ vanishes at the ‘vertical’ boundary of $T_{a}$. Extending these vector fields $X_{a}$ by zero outside $T_{a}$, we get a vector field $X$ on $M\times[0,1]$ representing the function $f$ on $A$. However, this vector field $X$ is clearly non-transverse to the boundaries $M\times\{0,1\}$ since it vanishes on an open set.
Remark 3.12.
The condition that the dynamical bordism $(W,X)$ representing a partial recursive function $f$ is clean is also of central relevance here. Intuitively it forces that the flow of $X$ captures topologically the operation of $f$. Indeed, since $(W,X)$ is constructed by gluing basic bordisms induced by a computable diffeomorphism, each of these basic tubes can only represent a simple intermediate calculation performed during the computation of $f$. Hence, locally the flow only executes simple operations and the involved dynamics of $f$ only emerges globally as a consequence of the topology of $W$, which determines how to combine these simple operations. In particular, the example of Remark 3.11 (and other variants) is non-clean either because $(M\times[0,1],X)$ is not constructed by gluing finitely many basic bordisms. This captures the fact that the flow of $X$ is very involved in this case, representing the whole calculation at once.
Remark 3.13.
The construction shown in Appendix A can be actually improved to allow an encoding that is robust under non-uniform perturbations. To be precise, a finite sequence $t=\{t_{i}\}$ can be encoded as the open set $U_{t}\subseteq I$ given by
@@
for $\varepsilon>0$ fixed and small. In other words, $U_{t}$ is the interval of length $(1-2\varepsilon)3^{-\ell-2}$ centered in $x_{t}$, where $x_{t}\in I$ is the point defined in Section 3.2. This induces a collection of open sets $\mathcal{U}_{\mathbb{N}}=\{U_{t}\}_{t\in\mathbb{N}}$ of $I$. By considering the collection $\mathcal{U}_{\Lambda}=\{U_{t^{+}}\times U_{t^{-}}\}$ for $t\in\Lambda$ a two-sided finite sequence with non-negative part $t^{+}$ and negative part $t^{-}$, we can also form a collection of open sets of $I^{2}$ encoding $\Lambda$. The whole construction of this paper can be carried out verbatim to work with this encoding in terms of open sets. Since the diameters of these open sets tend to zero as $\ell\to\infty$, we conclude that our construction is robust under non-uniform perturbations.
Furthermore, from the construction that we give in the Appendix, and using the fact that a partial recursive function can be implemented by different Turing machines, it is straightforward to prove the following fact.
Corollary 3.14.
The assignment $Z_{0}$ is not injective.
We finish this section by proving that the converse of Theorem 3.9 also holds, which completes the proof of Theorem 1.1. Explicitly, given any clean dynamical bordism $(W,X)$, the flow of $X$ defines a function $Z_{0}(W,X):(\mathbb{N}^{n})^{\sqcup m}\dashrightarrow(\mathbb{N}^{n})^{\sqcup m^{\prime}}$ that is computable.
@@
Now, suppose that the result is true for all clean bordisms formed by gluing $b\geq 1$ basic bordisms, and let us prove it for $b+1$. Let $W$ be a clean bordism whose underlying CW-complex $\widetilde{W}$ is formed by gluing $b+1$ basic bordisms and has reaching function $Z_{0}(\widetilde{W}):\mathbb{N}^{\sqcup m}\to\mathbb{N}^{\sqcup m^{\prime}}$. Choose one of the basic bordisms $M_{\varphi}=(M\times[0,1],X_{\varphi})$ in $\widetilde{W}$ and denote by $\widetilde{W}^{\prime}$ the result of removing $M_{\varphi}$ from $\widetilde{W}$. We have that $\widetilde{W}^{\prime}$ is composed of $b$ basic bordisms, so by induction hypothesis, the reaching function $Z_{0}(\widetilde{W}^{\prime})$ is computable.
First, suppose that the input and output boundaries of $M_{\varphi}$ are not boundaries of the bordism $\widetilde{W}$. Hence, removing $M_{\varphi}$ has the effect of adding a new input and output boundaries in $\widetilde{W}^{\prime}$, so that $Z_{0}(\widetilde{W}^{\prime}):\mathbb{N}^{\sqcup(m+1)}\to\mathbb{N}^{\sqcup(m^{\prime}+1)}$ and $Z_{0}(M_{\varphi}):\mathbb{N}\to\mathbb{N}$. For simplicity, let us assume that the newly created input and output boundaries correspond to the $(m+1)$-th copy of $\mathbb{N}^{\sqcup(m+1)}$ and the $(m^{\prime}+1)$-th copy of $\mathbb{N}^{\sqcup(m^{\prime}+1)}$, respectively. Let us denote by $(n,i)$ the point $n\in\mathbb{N}$ in the $i$-th component of $\mathbb{N}^{\sqcup m}$. In that situation, the function $Z_{0}(\widetilde{W})$ can be described in terms of $Z_{0}(\widetilde{W}^{\prime})$ and $Z_{0}(M_{\varphi})$ as follows
$Z_{0}(\widetilde{W})(n,i)=\left\{\begin{array}[]{ll}(n^{\prime},j)&\textrm{if }Z_{0}(\widetilde{W}^{\prime})(n,i)=(n^{\prime},j)\textrm{ with }j\leq m^{\prime},\\ Z_{0}(\widetilde{W}^{\prime})(Z_{0}(M_{\varphi})(n^{\prime}),m+1)&\textrm{if }Z_{0}(\widetilde{W}^{\prime})(n,i)=(n^{\prime},m^{\prime}+1).\end{array}\right.$
Using the inductive construction of partial recursive functions, it is easy to check that this function can be written as a recursion on $Z_{0}(\widetilde{W}^{\prime})$ and thus it is a partial recursive function provided that $Z_{0}(\widetilde{W}^{\prime})$ is so. The case in which $M_{\varphi}$ connects with an output boundary of $\widetilde{W}$ is completely analogous by adjusting the number of input and output components of the boundary of $\widetilde{W}^{\prime}$, and results in a simpler expression for $Z_{0}(\widetilde{W})$. This proves the inductive step and therefore the theorem follows. ∎
This result, jointly with Theorem 3.9, actually prove that clean dynamical bordisms have the same computational capabilities as Turing machines.

### v2

@@ 2.2. Partial Recursive Functions
[…]
@@
In this language, Kleene proved [23] that the function $f_{M}:\mathbb{N}\dashrightarrow\mathbb{N}$ implemented by a binary Turing machine $M$ is a partial recursive function. Reciprocally, every partial recursive function can be implemented via a Turing machine. Notice that the correspondence of the previous result is not bijective as different Turing machines can implement the same partial recursive function.
@@ 2.3. Turing Machines as finite state machines
The finite state machine $\mathcal{G}_{M}$ associated to the Turing machine $M$ allows us to visualize the changes in the internal states of $M$ while computing. At the beginning of the computation, the system starts at the initial vertex $q_{0}$ and, at each step, the system transitions to a new vertex through the edge corresponding to the symbol read in the tape at that moment, writing the tape and moving left or right according to the label of the edge.
@@
[…]
This Turing machine is designed to read a string of 0’s and 1’s and to return $1$ if it finds an even number of 1’s and 0 otherwise. According to the previous description, the associated finite state machine is the one depicted in fig. 4. The halting states are highlighted with a double border, and the initial state is marked with an incoming edge from no node.
$q_{2}$$q_{0}$$q_{1}$$(\square,1,-1)\;$(1,0,+1)(1,0,+1)(1,0,+1)(1,0,+1)(,0,1)\;($\square$,0,-1)(0,0,+1)(0,0,+1)(0,0,1)(0,0,-1)
@@ 3.2. Cantor-based encodings
The standard ternary Cantor set is the set of real numbers $x\in\mathbb{R}$ of the form $x=\sum_{i=1}^{\infty}\epsilon_{i}3^{-i}$, with $\epsilon_{i}=0$ or $2$. This Cantor set can be used to encode natural numbers (through their binary expansion) in the real line or, even more interestingly, paths in a binary tree ($\epsilon_{i}=0$ means that we turn left in the $i$-th branch, and $\epsilon_{i}=2$ that we turn right).
Using these ideas, we can define a useful encoding of finite binary sequences with symbols from the alphabet $\mathcal{A}=\{0,1,\square\}$ into the closed interval $I=[-1,1]\subseteq\mathbb{R}$. For this purpose, let $t=\{t_{n}\}$ be a finite sequence with $t_{n}\in\mathcal{A}$ for $n\geq 0$ and $t_{m}=\square$ for $m$ large enough. We encode $t$ in $I$ as the point
@@
The tuple $(\mathbb{D},X,\iota)$ is a marked dynamical germ, where $\mathbb{D}$ is the standard $n$-dimensional disk of radius $2$, $\iota:I^{n}\hookrightarrow\mathbb{D}^{n}$ the usual inclusion and $X_{p}=(0,-1)$ for all $p\in\mathbb{D}^{n}$ is the normal vector field. We shall refer to this marked dynamical germ as the standard disc.
@@ 3.3. Partial functions from dynamical bordisms
We may have $\psi_{(W,X)}(c)=\bullet$ for two reasons: Either the flow starting at $x$ does not hit the outgoing boundary $M^{\prime}$ or it hits $M^{\prime}$ but not in the image of the Cantor set. as represented in Figure 5.
Figure 5. Different possibilities occurring in the definition of $\psi_{(W,X)}:C^{n}\dashrightarrow C^{n}$.
@@
The clumsy condition that the line $\{n\}\times\kappa(\mathbb{N}^{n-1})$ must be mapped to a line $\{n^{\prime}\}\times\kappa(\mathbb{N}^{n-1})$ should be understood as the fact that the last n1n-1 components of the initial point are just “auxiliary data” that do not affect the final result of the calculation. The need for these auxiliary data is very clear geometrically, since the reaching function of a smooth flow is injective but we want to represent also non-injective functions. To avoid this problem, we use the vertical direction to store auxiliary data that allows us to separate points.
The general case of a dynamical bordism $(W,X):(M,X,\bm{\iota})\to(M^{\prime},X^{\prime},\bm{\iota}^{\prime})$ between non-connected objects can be worked similarly component-wise to get a partial map
$\psi_{(W,X)}=\sqcup_{j}\psi_{(W,X),j}:(\mathbb{N}^{n})^{\sqcup m}\dashrightarrow(\mathbb{N}^{n})^{\sqcup m^{\prime}},$
where $m$ is the number of components of $M$ and $m^{\prime}$ the one of $M^{\prime}$. As in the previous case, this induces a partial function $Z_{0}(W,X):\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$. We have thus proven the following result.
@@
At this point, an important question arises: What partial functions are representable by a clean dynamical bordism? In other words, given a partial function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, does there exist a clean dynamical bordism $W_{f}$ such that $Z_{0}(W_{f})=f$? In this direction, the following theorem establishes the first part of Theorem A, whose proof is provided in Section 4.
@@
In general, the topology of the bordism representing a partial recursive function $f:\mathbb{N}\to\mathbb{N}$ will not be trivial, capturing the complexity of the function. In particular, the function $f$ can be represented by a basic morphism if and only if $f$ extends to a diffeomorphism of the disc, when seen $\mathbb{N}\hookrightarrow\mathbb{D}$ via $\kappa$ (recall that $\mathbb{D}$ is the standard disk of radius $2$). However, not all the partial recursive functions can be extended, such as the function that reverses the binary expansion of a natural number, i.e., $f\left(\sum_{i=0}^{n}a_{i}2^{i}\right)=\sum_{i=0}^{n}a_{n-i}2^{i}$. Indeed, the points in $\mathbb{D}$ associated to the sequences $x_{n}=2^{n}$ and $x^{\prime}_{n}=2^{n}+2^{n-1}$ converge to the same limit $(-1,0)\in\mathbb{D}$, but $f(x_{n})=1$ and $f(x_{n}^{\prime})=3$ for all $n$.
Recall that, by definition, the vector field of a dynamical bordism is transverse to the incoming and outgoing boundaries of the bordism. This condition is crucial in Theorem 3.9. If we drop it, it is easy to construct a vector field on a topologically trivial bordism representing any function, just by extending the flowlines on the Cantor set by zero. Indeed, if $A\subseteq M$ is any discrete subset of a compact manifold $M$ (possibly with boundary), given any function $f:A\to A$, we can easily construct a degenerate non-transverse dynamical bordism representing it. To do so, in $M\times[0,1]$, choose pairwise disjoint paths $\gamma_{a}$ connecting $(a,0)$ with $(f(a),1)$ for all $a\in A$. Since $A$ is discrete, these paths $\gamma_{a}$ can be thickened into pairwise disjoint closed tubes $T_{a}\subseteq M\times[0,1]$. We endow each of these tubes $T_{a}$ with a flow-box vector field $X_{a}$ in such a way that the flow starting at $(a,0)$ hits $(f(a),1)$ and such that $X_{a}$ vanishes at the ‘vertical’ boundary of $T_{a}$. Extending these vector fields $X_{a}$ by zero outside $T_{a}$, we get a vector field $X$ on $M\times[0,1]$ representing the function $f$ on $A$. However, this vector field $X$ is clearly non-transverse to the boundaries $M\times\{0,1\}$ since it vanishes on an open set.
In the same spirit, notice that the condition that the dynamical bordism $(W,X)$ representing a partial recursive function $f$ is clean is also of central relevance here. Intuitively it forces that the flow of $X$ captures topologically the operation of $f$. Indeed, since $(W,X)$ is constructed by gluing basic bordisms induced by a computable diffeomorphism, each of these basic tubes can only represent a simple intermediate calculation performed during the computation of $f$.
@@
The construction of Theorem A can be improved to allow an encoding that is robust under non-uniform perturbations. To be precise, instead of encoding a finite sequence $t=\{t_{i}\}$ by a point in the square $I$, it can be encoded by a slightly bigger open set $U_{t}\subseteq I$ given by
@@
for $\varepsilon>0$ fixed and small. In other words, $U_{t}$ is the interval of length $(1-2\varepsilon)3^{-\ell-2}$ centered in $x_{t}$, where $x_{t}\in I$ is the point defined in Section 3.2. This induces a collection of open sets $\mathcal{U}_{\mathbb{N}}=\{U_{t}\}_{t\in\mathbb{N}}$ of $I$. By considering the collection $\mathcal{U}_{\Lambda}=\{U_{t^{+}}\times U_{t^{-}}\}$ for $t\in\Lambda$ a two-sided finite sequence with non-negative part $t^{+}$ and negative part $t^{-}$, we can also form a collection of open sets of $I^{2}$ encoding $\Lambda$. The whole construction of this paper can be carried out verbatim to work with this encoding in terms of open sets.
We finish this section by proving that the converse of Theorem 3.9 also holds, which completes the proof of Theorem A. Explicitly, given any clean dynamical bordism $(W,X)$, the flow of $X$ defines a function $Z_{0}(W,X):(\mathbb{N}^{n})^{\sqcup m}\dashrightarrow(\mathbb{N}^{n})^{\sqcup m^{\prime}}$ that is computable.
@@
[…]